OpenStudy (anonymous):

given the linear equation 2y+6=5x,find the y-intercept

OpenStudy (anonymous):

first , isolate y on one side

OpenStudy (anonymous):

that is , take everything else to the other side

OpenStudy (anonymous):

tell me what you get then

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what did you get ?

OpenStudy (anonymous):

trying to figure it out im not very good at linear equa tions

OpenStudy (anonymous):

is the answer 3

OpenStudy (jdoe0001):

2y+6=5x <-- subtract "6" from both sides, then divide by 2 both sides

OpenStudy (anonymous):

the possible answers are -3,3,-6/5,5/2

OpenStudy (anonymous):

no

OpenStudy (anonymous):

first , subtract 6 from both sided

OpenStudy (anonymous):

tell me what u get

OpenStudy (anonymous):

-4y+0

OpenStudy (anonymous):

NO 2y + 6 = 5x subtract 6 2y +6 - 6 = 5x - 6 2y = 5x - 6 divie by 2 2y/2 = 5x - 6/2 y = 5/2x - 3 this is of the form y = mx +c , where m is slope and c is y intercept , now what is the answer